# ProFL: Performative Robust Optimal Federated Learning

### v1

@@ 2 Problem Formulation
and
$\nabla\mathcal{L}_{i,2}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}\left[\ell(Z_{i};\theta)\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}\right].$
Note that PFL proposed by Jin et al., (2024) only updates the model with $\nabla\mathcal{L}_{i,1}(\theta)$ while completely neglecting $\nabla\mathcal{L}_{i,2}(\theta)$. This error can accumulate, and eventually, PFL can converge to a PS point far from the optimal. Thus, computing the PG for a more accurate estimate is crucial, but estimating $\nabla\mathcal{L}_{i,2}(\theta)$ is challenging. Although Izzo et al., (2021) proposed a method to estimate the PG, its performance is only evaluated in a centralized setting under a specific performative distribution map $D_{i}(\theta)$ (i.e., when $f_{i}(\theta)$ is a linear function of $\theta$) without considering data contamination. Moreover, it assumes that PR is convex, which is often unrealistic.
In summary, an FL algorithm that yields the PO point, even under a clean distribution and with rather limiting assumptions, has not been developed previously. This paper fills that gap by proposing an FL algorithm that provably converges to the PO point, even with non-convex PR, and handles data contamination.
Remark 2.1.
3 A Naive Extension from CS to FL
Consider first directly extending Izzo et al., (2021) to the FL setting. We shall call this naive algortihm Performative optimal Federated Learning (PoFL). The pseudocode is provided in Appendix A. In this algorithm, each active client $i$ in set $\mathcal{I}_{t}\subseteq\mathcal{V}$, after synchronization with the current global model $\overline{\theta}^{t}$, uses its local data to update the model from $\theta_{i}^{t}$ to $\theta_{i}^{t+R}$. During these updates, the data distribution changes, and consequently, client $i$ would use samples from the contaminated distribution $P_{i}(\theta_{i}^{t})$ to update $\theta_{i}^{t}$. To find the optimal ${\theta}^{\text{PO}}=\arg\min_{\theta}\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$, each client $i$ should estimate its local performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$ to update the local model $\theta_{i}^{t}$, i.e.,
@@ 3 A Naive Extension from CS to FL
By Eq. (3), $\nabla\mathcal{L}_{i}(\theta_{i}^{t})=\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ consists of two terms. PoFL estimates $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$ by simply averaging $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$. We denote the empirical risk on data drawn from distribution $Q_{i}$ or $D_{i}(\mathbf{\theta}_{i}^{t})$ as ^$\widehat{\nabla}\mathcal{L}_{i,1}(Q_{i},\theta_{i}^{t})$ and ^$\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$, respectively. Then the empirical risk on $P_{i}(\mathbf{\theta}_{i}^{t})$ is ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t}).$ Let ^$\widehat{\mathcal{L}}_{i,1}(\theta_{i}^{t})$ denotes the average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$. We denote ^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$, ^$\widehat{\nabla}\mathcal{L}_{i,2}(Q_{i},\theta_{i}^{t})$, and ^$\widehat{\nabla}\mathcal{L}_{i,2}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$ similarly. ^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ is estimated by a weighted average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$ with weight vector
$w(z_{j};\theta_{i}^{t}):=\left.\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p(z_{j};f_{i}(\theta))}{\partial f_{i}(\theta)}\right|_{\theta=\mathbf{\theta}_{i}^{t}}\in\mathbb{R}^{d}$ (5)
@@
Because the distribution map $D_{i}(\theta)=p(z;f_{i}(\theta))$ and $f_{i}(\theta)$ are unknown, we need to estimate $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}.$ Following Izzo et al., (2021), we adopt finite difference approximation $\left.\frac{df_{i}(\theta)}{d\theta}\right|_{\theta=\theta^{t}_{i}}\approx\Delta f_{i}(\Delta\theta_{i})^{\dagger}$. Specifically, at iteration $t$, each client $i$ collects the $H$ most recent models (excluding $\theta^{t}_{i}$) and forms a matrix $\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}$, then
$\Delta\theta_{i}$ $=\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}-\theta^{t}_{i}\textbf{1}^{T}_{H};$ (6)
$\Delta f_{i}$ $=\begin{bmatrix}f_{i}(\theta^{t-H}_{i})&\cdots&f_{i}(\theta^{t-1}_{i})\end{bmatrix}-f_{i}(\theta^{t}_{i})\textbf{1}^{T}_{H},$
@@
Where $\textbf{1}_{H}$ is an $H$-dimensional all-ones vector, both $\Delta f_{i}$ and $\Delta\theta_{i}$ are $d\times H$ matrices, and $(\Delta\theta_{i})^{\dagger}$ is the pseudo-inverse of $\Delta\theta_{i}$. For each iteration of each client, we know $\theta_{i}^{t}$, but do not know the exact value of $f_{i}(\theta_{i}^{t})$. Denote the estimator as $\widehat{f}_{i}(\mathcal{S}_{i})$, meaning we use the dataset $\mathcal{S}_{i}$ and the function $\widehat{f}_{i}(.)$ to estimate the value of $f_{i}(\theta_{i}^{t})$. This estimator is then used to compute an estimate of $\frac{df_{i}}{d\theta}$. The estimator of $\frac{df_{i}}{d\theta}$ is denoted as $\frac{\widehat{df_{i}}}{d\theta}$ and can be computed as $\frac{\widehat{df_{i}}}{d\theta}\approx\widehat{f}_{i}(\Delta\theta_{i})^{\dagger}$. Finally, $\frac{\widehat{df_{i}}}{d\theta}$ will be used to calculate ^$\widehat{\nabla}\mathcal{L}_{i,2}$. By combining this with ^$\widehat{\nabla}\mathcal{L}_{i,1}$, we obtain the performative gradient, allowing each client to update the local model $\theta_{i}^{t}$ using Eq. (4).
Concerns with POFL. Although Izzo et al., (2021) showed that the above method for estimating $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}$ was effective in centralized settings, it was evaluated under specific scenarios where $f_{i}(\theta)$ is linear in $\theta$ and data samples perfectly represent $D_{i}(\theta)$ without contamination. In practical FL settings with model-induced distribution shift, this method will perform poorly due to the following reasons:
  1. 1.
    Non-linearity of $f_{i}(\theta)$. When $f_{i}(\theta)$ is non-linear, ignoring the higher-order terms in the Taylor expansion of $\frac{\widehat{df_{i}}}{d\theta}$ may be biased. While Izzo et al., (2021) demonstrated that this method works well for linear $f_{i}(\theta)$, the impact of non-linearity on performance remains unclear.
  2. 2.
    Amplified sampling error. Since $\frac{df_{i}(\theta)}{d\theta}$ is estimated from limited samples, the sampling error of the estimation $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$ is inherited when computing PG, $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$. As shown in Lemmas 5.7, 5.8 and  5.9 in Section 5, the sampling error in estimating $\frac{df_{i}(\theta)}{d\theta}$ is amplified when computing $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$, making the performative gradient and algorithm more sensitive to sampling errors.
  3. 3.
    Contaminated data. Due to contamination $Q_{i}$, the data used for estimating $\frac{df(\theta)}{d\theta}$ and $f_{i}(\theta)$ does not follow the actual $D_{i}(\theta)$. Even a small error in a single model update for one client can cascade, affecting subsequent data and the entire network.
4 The Proposed Algorithm: PROFL
The previous section summarized the drawbacks of a straightforward extension of the centralized solution to the federated setting. To address all these weaknesses, we propose several simple yet effective strategies that lead to a robust new FL algorithm. We shall call it Performative robust optimal Federated Learning (ProFL). The pseudocode is provided in Algorithm 1. Below are its key features.
@@ 4 The Proposed Algorithm: PROFL
1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$, error bound $\Phi_{i}$ for all $i\in\mathcal{V}$.
2:  for $t=0$ to $T-1$ do
@@
7:           ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right).$
@@
11:           ^$\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right)$
@@
13:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
@@
15:           Input={^$=\{\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});H;d;\eta;\varphi_{i};\Delta f_{i}(\Delta\theta_{i})^{\dagger};\Phi_{i}\}$
@@
17:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
18:  Output: $\overline{\theta}^{T}$
  1. 1.
@@
    Handling complex distribution shifts with non-linear $f_{i}(\theta)$. ProFL employs the following strategies: (i) reducing the learning rate $\eta$ and the estimation window $H$; (ii) estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server. Reducing $\eta$ and $H$ helps reduce the estimation error resulting from non-linear $f_{i}(\theta)$ and stabilizes the algorithm. When the local sample size $n_{i}$ is small, $\widehat{f}_{i}(\mathcal{S}_{i})$ may significantly deviate from $f_{i}(\theta)$, leading to poor estimation of $\frac{df_{i}}{d\theta}$. Estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server mitigates the negative impact of sampling error.
  2. 2.
@@
    Balancing the impact of sampling error and computational costs. ProFL adaptively selects the sample size $n_{i}$ based on the error bounds of the PG and local loss. While a larger $n_{i}$ generally leads to a more accurate estimate, it may not always be feasible in FL due to clients’ limited computational capabilities. By using predefined error bounds and relevant computed values—such as model variance, loss function, gradient information, and learning rates—we can determine the lower bound for each client’s $n_{i}$, balancing sampling error and computational costs. As these values change during updates, $n_{i}$ should be selected adaptively. More details are provided in Section 5.3.
  3. 3.
    Handling data contamination. ProFL mitigates contamination $Q_{i}$ by reducing $\epsilon_{i}$ through local outlier identification and removal. Since the samples $\mathcal{S}_{i}\sim P_{i}(\theta_{i}^{t})$are noisy, and the raw data is not visible to the central server, contamination must be identified and cleaned locally before estimating the performative gradient $\nabla\mathcal{L}_{i}=\nabla\mathcal{L}_{i,1}+\nabla\mathcal{L}_{i,2}$. Given the contaminated dataset $\mathcal{S}_{i}$, each client in ProFL uses algorithm Robust gradient to identify “clean” samples $\mathcal{S}^{\prime}_{i}=\{z_{j}\in\mathcal{S}_{i}:z_{j}\sim D_{i}(\theta_{i}^{t})\}$ and compute robust gradient $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$. $\mathcal{S}^{\prime}_{i}$ are then used to estimate $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$. The outlier identification and removal mechanism reduces $\epsilon_{i}$ (the fraction of contaminated data) and can be selected based on prior knowledge of $\epsilon_{i},Q_{i}$ (if there are any). In Appendix A, we provide an example of Robust Gradient for a general case with unknown $\epsilon_{i}$ and arbitrary $Q_{i}$. It uses singular value decomposition (SVD) to detect gradients that meet two key properties: 1) having a significant impact on the model and potentially disrupting training (i.e., gradients with large magnitudes that systematically point in a specific direction); and 2) being far from the average in the vector space (i.e., “long tail” data). We iteratively remove gradients whose distance from the average exceeds a threshold. More details on the Robust Gradient mechanism and its comparison to prior work are provided in Appendix  A.
In section 5, we analyze both algorithms’ convergence and the impact of the factors above on performance. Compared to PFL Jin et al., (2024) and methods in Izzo et al., (2021); Miller et al., (2021), we relax assumptions and show convergence under weaker conditions.

### v2

@@ 3.2.1 Federated Estimation of the Performative Gradient
$\nabla\mathcal{L}_{i,2}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}\left[\ell(Z_{i};\theta)\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}\right]$}
Note that PFL jin2023performative only updates the model using $\nabla\mathcal{L}_{i,1}(\theta)$ and neglects $\nabla\mathcal{L}_{i,2}(\theta)$. As a result, estimation error accumulates, and the algorithm may converge to a PS point that is far from PO point. Therefore, accurately estimating the PG is critical; however, estimating $\nabla\mathcal{L}_{i,2}(\theta)$ remains a significant challenge in FL, particularly under data contamination.
In ProFL, each active client $i$ in set $\mathcal{I}_{t}\subseteq\mathcal{V}$, after synchronization with the current global model $\overline{\theta}^{t}$, uses its local data to update the model from $\theta_{i}^{t}$ to $\theta_{i}^{t+R}$. During these updates, the data distribution evolves with the model, and consequently, client $i$ uses a sample set $\mathcal{S}_{i}$ drawn from the contaminated distribution $P_{i}(\theta_{i}^{t})$ to update its local model $\theta_{i}^{t}$. To find the optimal ${\theta}^{\text{PO}}=\arg\min_{\theta}\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$, each client $i$ should estimate its local performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$ to update the local model $\theta_{i}^{t}$, i.e.,
@@
By Eq. (4), $\nabla\mathcal{L}_{i}(\theta_{i}^{t})=\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ consists of two terms. To estimate $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$, ProFL applies a robust gradient strategy that filters the contaminated sample set. Specifically, a refined subset $\mathcal{S}^{\prime}_{i}$ is first extracted from the original sample set $\mathcal{S}_{i}\sim P_{i}(\theta_{i}^{t})$ via
$\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}(\mathcal{S}_{i};\theta^{t}_{i};J_{1};J_{2}),$
where $J_{1}$ and $J_{2}$ are hyperparameters controlling the selection criteria. The final estimate $\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})$ is obtained by averaging $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples $z_{j}\in\mathcal{S}^{\prime}_{i}$, where $\mathcal{S}^{\prime}_{i}$ is the cleaned sample set selected via the robust gradient strategy (detailed in Section 3.3).
For reference, we denote the empirical gradient estimate on samples drawn from distributions $Q_{i}$ and $D_{i}(\theta_{i}^{t})$ as $\widehat{\nabla}\mathcal{L}_{i,1}(Q_{i},\theta_{i}^{t})$ and $\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$, respectively. Similarly, the empirical risk over cleaned samples is denoted as $\widehat{\mathcal{L}}_{i,1}(\theta_{i}^{t})$, computed by averaging $\ell(z_{j};\theta_{i}^{t})$ over $z_{j}\in\mathcal{S}^{\prime}_{i}$.
We also define $\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ and its counterparts $\widehat{\nabla}\mathcal{L}_{i,2}(Q_{i},\theta_{i}^{t})$ and $\widehat{\nabla}\mathcal{L}_{i,2}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$ analogously. Specifically, $\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ is computed as a weighted average of $\ell(z_{j};\theta_{i}^{t})$ over the cleaned sample set $\mathcal{S}^{\prime}_{i}$, with weights determined by sensitivity to model-induced distributional parameters. The weight vector is defined as
$w(z_{j};\theta_{i}^{t}):=\left.\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p(z_{j};f_{i}(\theta))}{\partial f_{i}(\theta)}\right|_{\theta=\mathbf{\theta}_{i}^{t}}\in\mathbb{R}^{d}$}, (6)
@@
which encodes how the density $p(z_{j};f_{i}(\theta))$ changes with respect to the model parameter $\theta$ through the mapping $f_{i}(\theta)$. Since the performative distribution $D_{i}(\theta)=p(z;f_{i}(\theta))$ and $f_{i}(\theta)$ are both unknown in practice, we must estimate both $f_{i}(\theta)$ and its Jacobian $\frac{df_{i}(\theta)}{d\theta}$ from the observed data. Following pmlr-v139-izzo21a, we adopt finite difference approximation $\left.\frac{df_{i}(\theta)}{d\theta}\right|_{\theta=\theta^{t}_{i}}\approx\Delta f_{i}(\Delta\theta_{i})^{\dagger}$. Specifically, at iteration $t$, each client $i$ collects the $H$ most recent models (excluding $\theta^{t}_{i}$) and forms a matrix $\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}$, then
$\Delta\theta_{i}=\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}-\theta^{t}_{i}\textbf{1}^{T}_{H};\quad\Delta f_{i}$ $=\begin{bmatrix}f_{i}(\theta^{t-H}_{i})&\cdots&f_{i}(\theta^{t-1}_{i})\end{bmatrix}-f_{i}(\theta^{t}_{i})\textbf{1}^{T}_{H},$
@@
Where $\textbf{1}_{H}$ is an $H$-dimensional all-ones vector, both $\Delta f_{i}$ and $\Delta\theta_{i}$ are $d\times H$ matrices, and $(\Delta\theta_{i})^{\dagger}$ is the pseudo-inverse of $\Delta\theta_{i}$. For each iteration of each client, we know $\theta_{i}^{t}$, but do not know the exact value of $f_{i}(\theta_{i}^{t})$. Denote the estimator as $\widehat{f}_{i}(\mathcal{S}^{\prime}_{i})$, meaning we use the dataset $\mathcal{S}^{\prime}_{i}$ and the function $\widehat{f}_{i}(.)$ to estimate the value of $f_{i}(\theta_{i}^{t})$. This estimator is then used to compute an estimate of $\frac{df_{i}}{d\theta}$. The estimator of $\frac{df_{i}}{d\theta}$ is denoted as $\frac{\widehat{df_{i}}}{d\theta}$ and can be computed as $\frac{\widehat{df_{i}}}{d\theta}\approx\widehat{f}_{i}(\Delta\theta_{i})^{\dagger}$. Finally, $\frac{\widehat{df_{i}}}{d\theta}$ will be used to calculate $\widehat{\nabla}\mathcal{L}_{i,2}$. By combining this with $\widehat{\nabla}\mathcal{L}_{i,1}$, we obtain the performative gradient, allowing each client to update the local model $\theta_{i}^{t}$ using Eq. (5).
@@
1:Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$, error bound $\Phi_{i}$ for all $i\in\mathcal{V}$, threshold $J_{1},J_{2}\in(0,1)$, number of clusters $C$.
2:for $t=0$ to T1T-1 do
@@
7:    $\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i};J_{1};J_{2}\right).$
@@
11:     $\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right)$
@@
13:    Compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$ by $\Delta\theta_{i}$ and $\Delta f_{i}$.
@@
[…]
@@
17:    Lines 18–20 are necessary only for small $n_{i}$.
18:    Send $\widehat{\frac{df_{i}}{d\theta}}$ to the server.
19:    $\widehat{\frac{df_{c}}{d\theta}}\leftarrow$ ServerAggregation $(\widehat{\frac{df_{i}}{d\theta}},C)$
20:    Compute $\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})$ using the aggregated $\widehat{\frac{df_{c}}{d\theta}}$.
@@
21:  If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
22:Output: $\overline{\theta}^{T}$
23:Server Aggregation:
24: Use $k$-means clustering to group clients into $C$ clusters.
25: Compute the aggregated gradient estimate $\widehat{\frac{df_{c}}{d\theta}}$ by averaging $\widehat{\frac{df_{i}}{d\theta}}$ for all $i$ in cluster $c$, for each cluster $c\in\{1,\ldots,C\}$.
26: Send $\widehat{\frac{df_{c}}{d\theta}}$ back to the clients belonging to cluster $c$ for all $c\in\{1,\ldots,C\}$.
3.3 Details of Robust Strategies
Although pmlr-v139-izzo21a demonstrated that the method for estimating $f_{i}(\theta)$ and its Jacobian $\frac{df_{i}(\theta)}{d\theta}$ is effective in centralized settings, their analysis relied on that $f_{i}(\theta)$ is linear in $\theta$ and the data perfectly reflects the clean distribution $D_{i}(\theta)$. In federated learning, however, model-induced distribution shifts are often nonlinear and local data may be limited or contaminated, which significantly undermines the accuracy of the PG estimation used in pmlr-v139-izzo21a. To overcome these challenges, ProFL incorporates a set of carefully designed robustness strategies, described in detail below.
@@ 3.3 Details of Robust Strategies
Non-linearity of $f_{i}(\theta)$. When $f_{i}(\theta)$ is non-linear, ignoring the higher-order terms in the Taylor expansion of $\frac{\widehat{df_{i}}}{d\theta}$ may be biased. While pmlr-v139-izzo21a demonstrated that this method works well for linear $f_{i}(\theta)$, the impact of non-linearity on performance remains unclear. To mitigate the error introduced by non-linearity, ProFL employs the following strategies: (i) reducing the learning rate $\eta$ and the estimation window $H$; (ii) estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server. Reducing $\eta$ and $H$ helps reduce the estimation error resulting from non-linear $f_{i}(\theta)$ and stabilizes the algorithm. When the local sample size $n_{i}$ is small, $\widehat{f}_{i}(\mathcal{S}_{i})$ may significantly deviate from $f_{i}(\theta)$, leading to poor estimation of $\frac{df_{i}}{d\theta}$. Estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server mitigates the negative impact of sampling error.
@@
Balancing the impact of sampling error and computational costs. In practice, $\frac{df_{i}(\theta)}{d\theta}$ must be estimated from finite samples, which introduces sampling error into the computation of the performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$. As shown in Lemmas 4.7, 4.8 and  4.9 in Section 4, the sampling error in estimating $\frac{df_{i}(\theta)}{d\theta}$ is amplified when computing $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$, making the performative gradient and algorithm more sensitive to sampling errors. To mitigate this, ProFL adaptively selects the sample size $n_{i}$ based on the error bounds of the PG and local loss. While a larger $n_{i}$ generally leads to a more accurate estimate, it may not always be feasible in FL due to clients’ limited computational capabilities. Therefore, ProFL balances estimation accuracy and resource constraints by computing a theoretical lower bound on $n_{i}$ using quantities such as model variance, loss curvature, learning rate, and empirical gradient variability. As these values change during updates, $n_{i}$ should be selected adaptively. More details of the analysis are provided in Section 4.3.
Handling data contamination. Due to contamination $Q_{i}$, the data used to estimating $\frac{df(\theta)}{d\theta}$ and $f_{i}(\theta)$ does not follow the actual $D_{i}(\theta)$. Even a small update error from one client can cascade and impact future data and the entire network. In FL, because data must remain local for privacy reasons, contamination must also be detected and removed locally rather than at the central server. Accordingly, ProFL reduce $\epsilon_{i}$ via local outlier identification and removal. To estimate $\nabla\mathcal{L}_{i,1}$, we should average $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $D_{i}({\theta}_{i}^{t})$; however, the samples actually collected by client $i$ are drawn from $P_{i}(\theta_{i}^{t})$and thus include an unknown fraction $\epsilon_{i}$ of outliers from an unknown distribution $Q_{i}$. ProFL needs to identify contaminated data and eliminate the corresponding contaminated gradients. We propose a mechanism that can remove these contaminated gradients and estimate gradients that are robust to outliers.
Importantly, ProFL is independent of the specific outlier identification and removal mechanism used and can function with any such mechanism. In our paper, we assume that contaminated data can follow an arbitrary distribution. In practice, additional information about the contamination process may be available based on the application or context, allowing us to design more efficient and less expensive outlier identification mechanisms leveraging this knowledge.
Algorithm 2 Robust Gradient
1:Input: Set $\mathcal{S}_{i}$ of samples $(z_{j})_{j=1}^{n_{i}}$, thresholds $J_{1},J_{2}\in(0,1)$, model $\theta^{t}_{i}\in\Theta$
2: Let $\widehat{\nabla}\leftarrow\frac{1}{|\mathcal{S}_{i}|}\sum_{z_{j}\in\mathcal{S}_{i}}\nabla\ell(z_{j};\theta^{t}_{i})$.
3: Let $\left[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right]_{z_{j}\in\mathcal{S}_{i}}$ be the $|\mathcal{S}_{i}|\times d$ matrix.
4: Apply SVD to this matrix and find top right singular vector $v$.
5: Compute $\tau_{j}=\left(\left(\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right)^{T}v\right)^{2}$ of $\nabla\ell(z_{j};\theta^{t}_{i})$.
6: Divide the interval $0,\max\tau_{j}]$ into $B$ equal-length segments. Let $\phi_{k}$ be the number of $\tau_{j}$ in the $k$-th segment.
7: Find the smallest $k\in\{1,\cdots,B\}$ with $\phi_{k}<J_{1}\cdot|\mathcal{S}_{i}|$. Set $\phi$ as the lower bound of the $k$-th segment.
9: Compare the average gradients $\widehat{\nabla}$ of $\mathcal{S}_{i}$ and $\widehat{\nabla}^{\prime}$ of $\mathcal{S}^{\prime}_{i}$. If $\frac{\|\widehat{\nabla}-\widehat{\nabla}^{\prime}\|}{\|\widehat{\nabla}\|}<J_{2}$ or $|\mathcal{S}^{\prime}_{i}|\leq\frac{n_{i}}{2}$, return $\mathcal{S}^{\prime}_{i}$ and $\widehat{\nabla}^{\prime}$; otherwise, set $\mathcal{S}_{i}\leftarrow\mathcal{S}^{\prime}_{i}$ and repeat from line 2.
As shown in Algorithm 2, our mechanism takes noisy samples $\mathcal{S}_{i}$ and current local model $\theta_{i}^{t}$ as inputs. The mechanism first computes the gradients $\nabla\ell(z_{j};\theta_{i}^{t})$ of all samples and iteratively identifies and removes the contaminated gradients. Since both the distribution $Q_{i}$ and fraction $\epsilon_{i}$ of contaminated data are unknown, we consider gradients contaminated if they exhibit two crucial properties: 1) they have large effects on the learned model and can disturb the training process (i.e., gradients have large magnitudes and systematically point in a specific direction); 2) they are located farther away from the average in the vector space (i.e., “long tail” data). To detect gradients that satisfy these properties, we adapt the approach in pmlr-v97-diakonikolas19a, which leverages singular value decomposition (SVD).
Specifically, we construct a matrix $\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}]_{z_{j}\in\mathcal{S}_{i}}$ using the centered gradients and find the top right singular vector $v$, where $\widehat{\nabla}$ represents the average of all gradients (lines 2-4). The centered gradients that are closer to $v$ are more likely to be outliers and we assign each gradient an outlier score $\tau_{j}$ as follows:
Given the outlier scores $\{\tau_{j}\}$, the gradients with scores above a threshold $\phi$ are considered contaminated and are discarded.
Key Innovations of Robust Gradient in ProFL. Unlike the method in pmlr-v97-diakonikolas19a, which requires a manually preselected threshold and resampling after each removal, ProFL determines the outlier threshold $\phi$ automatically from the empirical distribution of outlier scores $\tau_{j}$ and use the same dataset to get the roustabout gradient because it is difficult to determine without any prior knowledge of the data and sampled new data after each time of removal in impractical in PP.
Since we assume the outliers lie in the “long tail,” we divide the range $0,\max\tau_{j}]$ into $B$ equal-length segments and compute the relative frequency of each. The threshold $\phi$ is chosen as the smallest segment whose relative frequency falls below a predefined level. This method exploits the score concentration and tail sparsity properties, making the filtering process predictable and stable. The final robust gradient estimate $\widehat{\nabla}^{\prime}$ is computed by averaging gradients with scores below $\phi$, ensuring resilience to noise and contamination without requiring prior knowledge.